Improper Integral (Integrals with Infinite Limits)


∫dy/√(y-1) from 5 to ∞


1
Expert's answer
2021-05-07T11:40:32-0400

Solution:



"\int \frac{1}{\sqrt{y-1}}\,dx \quad from \quad 5\ to \ \infty"


can be written as


"\lim_{a \to \infty} \int_{5}^a \frac{1}{\sqrt{y-1}}\,dy"

"y-1 =t \implies\,dy=dt\\
y\rightarrow5; t\rightarrow4\ and\ y\rightarrow\infty; t\rightarrow\infty\\
\therefore \quad \lim_{a \to \infty} \int_{4}^a \frac{1}{\sqrt{t}}\,dt\\""\lim_{a \to \infty} \lbrack \lbrack \frac{t^{-\frac{1}{2}+1}}{-\frac{1}{2}+1}\rbrack _{a}-\lbrack \frac{t^{-\frac{1}{2}+1}}{-\frac{1}{2}+1}\rbrack _{4}\rbrack"

"\lim_{a \to \infty} \lbrack 2\sqrt{a} -2\sqrt{4}\rbrack \\
\lim_{a \to \infty} \lbrack 2\sqrt{a} -4\rbrack\\"

As a tends to infinite, entire expression tends to infinite.

Thus the given integral is divergent.